# Advances and Open Challenges in Federated Learning with Foundation Models

### v2

@@ VI-D FedFM Contribution Evaluation
TABLE VII: Summary of Evaluations of Traditional FL Models
@@ VI-D1 Contribution Evaluation for Traditional FL
The performance of the global FL model highly relies on the quality of the local dataset. Client and sample contribution evaluation can help the FL server analysis this through tracing back to local data, and is important to model aggregation. FL client and sample contribution evaluation methods can be divided into two main categories: 1) Shapley value (SV)-based evaluation and 2) influence-based evaluation (Table VII).
@@ VI-D3 Contribution Evaluation for FedFM
As for evaluating FedFM, so far, there have been few relevant works. The work [197] proposed a logic rule learning approach to select the optimal chain-of-thoughts prompts for improving the interpretability of federated prompt selection for multi-domain FMs. They cast this problem as a bi-level program, and solve it through variational expectation maximization. The work [198] proposed to evaluate FedFM by comparing the performance of the federated parameter-efficient fine-tuned model with traditional centralized fine-tuning methods. FedIT [199] leverages FL framework for FM instruction tuning and conducted studies on the widely used GPT-4 to exploit the heterogeneous and diverse sets of instructions. FedNLP [200] is a benchmark framework for evaluating FL methods on four common formulations of NLP tasks: text classification, sequence tagging, question answering and seq2seq generation. They proposed a universal interface between Transformer-based language models (e.g., BERT, BART) and FL methods under various non-IID partitioning strategies.
SV-based methods have been extensively employed in conventional FL scenarios to assess the contribution of each participant. However, applying SV-based methods to measure contributions in FedFM settings presents two notable challenges: inference latency and compositional gap.
@@ VI-E FedFM Reward Distribution
Similar to traditional FL, FedFM reward distribution needs to address the following issues.
@@ VI-E3 Cost Balancing
VII FedFM Frameworks
Setting up FedFM involves a multi-faceted approach involving both the software and hardware aspects to enable distributed training and fine-tuning across diverse devices while maintaining data privacy and security. It can be significantly facilitated by well-designed platforms and libraries supporting FL and the unique requirements of FMs. Here, we discuss leading FL frameworks which support large-scale FMs.
TABLE VIII: Comparative Analysis of FL Frameworks that Support FMs
Frameworks
No. of Supported
@@ 3 LLaMa (7B)
An overview of prospective platforms and libraries for FedFM are shown in Table VIII. We focus on frameworks such as FedML [204], FederatedScope-LLM [205], Fate-LLM [206], and OpenFedLLM[54], comparing the methods supported, dataset compatibility and maximum model sizes as showcased in their examples. The findings reveal a trend towards integrating state-of-the-art techniques to facilitate FedFM. However, there appears to be little emphasis on computational efficiency. Notably, all frameworks have been assessed with contemporary large-scale FMs, underscoring their practical relevance. Utilizing these frameworks can facilitate the reassessment of techniques previously surveyed with actual FMs, potentially catalyzing novel research directions.
@@
Implementing FedFM is an emerging and evolving research area, with continuously improving tools and practices. Selecting the right combination of platforms and libraries depends on the specific requirements of FedFM tasks, including the type of FMs involved, privacy, security and scalability requirements, as well as the computational resources available.
VIII Challenges and Promising Directions
@@
VIII-A Memory and Computing Challenges facing FedFM
@@
High Demand for Memory Resources: The large scale of FMs, often involving billions of parameters, results in significant memory demands. This is particularly challenging in FL environments, where client devices may have limited memory capacities. The key challenge is developing strategies to accommodate these large models efficiently within the memory constraints of diverse client devices.
High Demand for Computing Resources: In FedFM, the demand for high computational resources on FL client devices poses a persistent challenge. Despite strategies to mitigate this, it remains an open issue. Approaches such as model splitting [207] and model compression [208] have been proposed. In addition, efforts are made to develop lighter models and more efficient training algorithms [209] that are less taxing on device resources. Yet, these solutions only offer a partial remedy. The research on federated optimization [210, 25] indicate that finding the right balance between maintaining model performance and managing resource constraints is an ongoing struggle. Although these advances mark progress, ensuring that FedFM is adaptable for devices with diverse capabilities continues to be a significant open research question.
@@
VIII-B Communication Challenges facing FedFM
@@
Security Issues: In FedFM supported by advanced network technologies like 5G and beyond (B5G), security issues pose a significant challenge. The diversity in computing and communication capabilities across different participants in the network, stemming from variations in hardware (e.g., CPU, GPU), network connections (e.g., 4G, 5G, B5G, WiFi) and energy resources (e.g., battery, charged), leads to system heterogeneity  [236, 237, 238]. This diversity can introduce inconsistencies and vulnerabilities in FedFM [239]. Moreover, the presence of unreliable devices within the network might lead to Byzantine failures as highlighted in [240, 241]. These failures refer to scenarios where certain nodes in the network act in a faulty or malicious manner, further complicating the security landscape of FedFM. The varying levels of security across different devices exacerbate the difficulty in defending against attacks and ensuring system reliability. Despite ongoing efforts to enhance FL security [28], these inherent vulnerabilities in FedFM remain open.
@@
VIII-C Promising Direction: Quantum Computing for FedFM
@@
The ability of quantum computing to efficiently perform certain types of sampling and approximation (e.g., Monte Carlo sampling) can be advantageous for many ML tasks. For example, [250] have demonstrated a framework for reservoir computing with nonlinear quantum reservoirs, showcasing the computational capabilities across classical and quantum regimes. Quantum systems enable more efficient data structure representation, which can expedite computation in NNs during inference. When fine-tuning FedFM models for specific tasks, feature selection is a crucial process. Quantum computing can offer efficient solutions for this process, enhancing the performance of the resulting FMs [251]. Lastly, quantum parallelism can support simultaneous evaluation of different hyper-parameter settings. This capability is significant for the hyper-parameter tuning process, a time-consuming aspect of FM fine-tuning. Moreover, the capability of quantum computing for handling high-dimensional spaces can aid in optimizing the high-dimensional parameter space of FMs more efficiently.
@@
The integration of quantum and classical optimization steps in hybrid algorithms offers a balanced approach. By utilizing quantum processors for the more computationally demanding parts and classical processors for other parts, these hybrid models leverage the strengths of both quantum and classical computing. They can be particularly effective in overcoming current limitations of quantum computing, while still harnessing its advantages for optimization tasks [255]. Global optimization is a key factor in FedFM settings, where the objective is to achieve consensus among distributed models. Quantum algorithms can efficiently aggregate local updates and conduct global optimization. This can produce more effective and cohesive FMs in a collaborative learning environment. The incorporation of quantum optimization into FedFM represents a promising direction. By enhancing the efficiency of finding optimal solutions, these quantum-based techniques have the potential to significantly improve the performance and effectiveness of FedFM models.
@@
VIII-D Promising Direction: Evaluation of FedFM

### v4

@@ VI-D FedFM Contribution Evaluation
TABLE VI: Summary of Evaluations of Traditional FL Models
@@ VI-D1 Contribution Evaluation for Traditional FL
The performance of the global FL model highly relies on the quality of the local dataset. Client and sample contribution evaluation can help the FL server analysis this through tracing back to local data, and is important to model aggregation. FL client and sample contribution evaluation methods can be divided into two main categories: 1) Shapley value (SV)-based evaluation and 2) influence-based evaluation (Table VI).
@@ VI-D3 Contribution Evaluation for FedFM
As for evaluating FedFM, so far, there have been few relevant works. The work [233] proposed a logic rule learning approach to select the optimal chain-of-thoughts prompts for improving the interpretability of federated prompt selection for multi-domain FMs. They cast this problem as a bi-level program, and solve it through variational expectation maximization. The work [234] proposed to evaluate FedFM by comparing the performance of the federated parameter-efficient fine-tuned model with traditional centralized fine-tuning methods. FedIT [235] leverages FL framework for FM instruction tuning and conducted studies on the widely used GPT-4 to exploit the heterogeneous and diverse sets of instructions. FedNLP [236] is a benchmark framework for evaluating FL methods on four common formulations of NLP tasks: text classification, sequence tagging, question answering and seq2seq generation. They proposed a universal interface between Transformer-based language models (e.g., BERT, BART) and FL methods under various non-IID partitioning strategies. SV-based methods have been extensively employed in conventional FL scenarios to assess the contribution of each participant. However, applying SV-based methods to measure contributions in FedFM settings presents two notable challenges: inference latency and compositional gap.
@@ VI-E FedFM Reward Distribution
Similar to traditional FL, FedFM reward distribution, shown in Fig. 13, needs to address the following issues.
Figure 13: Illustration of reward distribution in FedFM.
@@ VI-E3 Cost Balancing
VII FedFM Implementation and Applications
Implementing FedFM requires careful consideration of both hardware and platform capabilities for handling the computational and storage demands of large-scale models. Here, we provide a concise tutorial on the key components and tools commonly used in FedFM applications, with reference to platforms in Table VII and representative applications in Table VIII.
VII-A FedFM Implementation
Setting up FedFM requires a multi-faceted approach involving both software and hardware aspects to enable distributed training and fine-tuning across diverse devices, while maintaining data privacy. It can be facilitated by well-designed platforms and libraries supporting FL and the requirements of FMs. Here, we discuss leading FL frameworks which support large-scale FMs, including advanced hardware configurations and specialized platforms.
TABLE VII: Comparative Analysis of FL Platforms and Libraries that Support FMs.
Platforms
No. of
@@ 3 LLaMa (7B)
VII-A1 Platform and Library Considerations
An overview of prospective platforms and libraries for FedFM are shown in Table VII. We focus on frameworks including FedML [240], FederatedScope-LLM [241], FATE-LLM [242], and OpenFedLLM [72], comparing the methods supported, dataset compatibility and maximum model sizes as showcased in their examples.
  • FedML [240] is a flexible and scalable platform designed for implementing and experimenting with various FL models. It supports both simulation and real-world deployment, making it an ideal choice for research and practical applications of FedFM. FedML’s integration with popular machine learning frameworks like TensorFlow and PyTorch allows for seamless adaptation of foundation models in federated settings.
  • FederatedScope-LLM [241] is an extension of FederatedScope focused on LLMs. It addresses the unique challenges of training LLMs in a federated environment, such as efficient communication and privacy preservation. This platform is relevant for FedFM as it provides specialized tools to manage the complexities of FedFM.
  • FATE-LLM [242], a specialized version of the Federated AI Technology Enabler (FATE) platform, is tailored for LLMs. It leverages secure computation protocols and privacy-preserving techniques, making it suitable for FedFM applications that require strong data security and compliance with privacy regulations.
  • OpenFedLLM [72] is an open-source framework aimed at facilitating the FedFM. Its extensibility and ease of use make it a valuable tool for developing FedFM workflows. OpenFedLLM supports various FL paradigms and integrates with existing ML libraries, promoting innovation in FedFM.
@@ VII-A1 Platform and Library Considerations
The findings reveal a trend of integrating state-of-the-art techniques to facilitate FedFM. However, there appears to be little emphasis on computational efficiency. Notably, all frameworks have been assessed with contemporary large-scale FMs, underscoring their practical relevance. Utilizing these frameworks can facilitate the reassessment of techniques previously surveyed with actual FMs, potentially inspiring novel research directions. Implementing FedFM is an emerging and evolving research area, with continuously improving tools and practices. Selecting the right combination of platforms and libraries depends on the specific requirements of FedFM tasks, including the type of FMs involved, privacy, security and scalability requirements, as well as the computational resources available.
VII-A2 Hardware Considerations
High-performance central servers are crucial for coordinating the FedFM process. These servers, often hosted on cloud platforms such as AWS [^2] https://aws.amazon.com/, Google Cloud [^3] https://cloud.google.com/?hl=en, Microsoft Azure [^4] https://azure.microsoft.com/en-gb or NVIDIA FLARE [^5] https://developer.nvidia.com/flare, need to have high-end configurations including multi-GPU setups, large memory, and extensive storage solutions to manage the aggregation and training of large-scale models. Distributed edge servers, equipped with multiple GPUs, high-core-count CPUs, and large memory, act as intermediaries between the central server and edge devices to reduce latency and enhance scalability.
Regarding edge devices, high-performance smartphones and tablets equipped with the latest processors, substantial RAM, and storage capabilities are suitable for initial phases of model training and inference tasks. Powerful laptops and desktops with multi-core CPUs, ample RAM, and dedicated GPUs are essential for edge clients participating in FedFM. In addition, advanced IoT devices with enhanced processing units, such as NVIDIA Jetson series [^6] https://www.nvidia.com/en-us/autonomous-machines/embedded-systems/ or Google Coral AI accelerators, can support model inference and lightweight training tasks in distributed settings.
VII-A3 Practical Implementation
For simulation and testing, high-performance computing clusters equipped with multiple GPUs and high-speed interconnects can mimic large-scale federated environments and provide insights into performance bottlenecks. Cloud-based simulation platforms such as AWS EC2 with p3/p4 instances, Google Cloud with TPU support, and Microsoft Azure with NDv2/NDv4 instances offer scalable simulation and testing environments for FedFM setups.
Deployment of FedFM systems benefits from containerization and orchestration tools like Docker and Kubernetes [^7] https://kubernetes.io/. Kubernetes operators can manage the lifecycle of FL jobs, ensuring efficient resource allocation and fault tolerance. For edge deployment, tools like NVIDIA Fleet Command [^8] https://resources.nvidia.com/en-us-retail-and-edge-ai/fleet-command-web-page?xs=205038&ncid=no-ncid and AWS IoT Greengrass [^9] https://aws.amazon.com/greengrass/ facilitate the deployment and management of FL models on edge devices, supporting secure and efficient model inference and updates in distributed environments.
VII-B FedFM Applications
For NLP applications, FedFM enables collaborative training of FMs across diverse datasets, enhancing language understanding and generation capabilities while preserving data privacy. For speech recognition, FedFM facilitates the training of robust models that can generalize across different accents and languages, leveraging distributed data without centralizing sensitive voice data. In recommender systems, FedFM supports the development of personalized recommendations by aggregating user preferences from multiple sources, ensuring privacy and enhancing recommendation accuracy. In healthcare, FedFM allows for the aggregation of medical data from various institutions, enabling the development of more accurate diagnostic models and personalized treatment plans without compromising patient privacy. The FedFM applications are shown in Table VIII and the detailed descriptions of each application are given as follows.
TABLE VIII: Related Works on the Applications of FedFM
Application (Modality) Related Works Year Task and Domain On-Device Personalization
Healthcare (Vision and Text) FedPR [243] 2023 MRI Reconstruction No No
FedTherapist [244] 2023 Mental Health Prediction Yes No
Recommender System (Text) GPT-FedRec [245] 2024 General No No
TransFR [246] 2024 General Yes Yes
PPLR [247] 2024 General No Yes
Speech (Audio) FedE2EASR [248] 2023 Speech Recognition No Yes
FedASR [249] 2023 Speech Recognition Yes Yes
pFedS2T [250] 2024 Speech-to-Text No Yes
Multilingual NLP (Text) MFPT [251] 2024 Multi-Tasks Yes No
FL-MetaSend [252] 2024 Machine Translation No No
Fed-MNMT [253] 2023 Machine Translation No No
PMMFL [254] 2022 Multi-Tasks No No
FedKC [255] 2022 Language Understanding No No
  • In healthcare, FedPR [243] focused on magnetic resonance imaging (MRI), using an FM pre-trained on public datasets and training visual prompts from decentralized clinical data via a personalized FL mechanism, achieving competitive performance with limited local data. Shin et al.[244] developed a mobile mental health monitoring system that uses user speech and keyboard input to fine-tune FedFM, demonstrating high accuracy in predicting mental health conditions such as depression, stress, and mood.
  • For recommender systems, GPT-FedRec [245] utilized ChatGPT’s strong zero-shot generalization ability in federated recommendation, facilitating hybrid retrieval and re-ranking of recommendation systems. Zhang et al. [246] introduced a transparent federated recommendation framework that enhances recommendation accuracy while maintaining user privacy. Zhao et al. [247] presented a personalized privacy-preserving language representation learning method, using FL to protect user data privacy.
  • In the speech domain, Azam et al. [248] proposed federated end-to-end automatic speech recognition systems, leveraging FedFM’s advantages in privacy-sensitive audio scenarios. FedASR [249] showed that FL can improve automatic speech recognition by training collaboratively on decentralized audio data. Du et al. [250] introduced personalized federated speech-to-text models, addressing the need for accurate speaker-independent performance while offering personalized adaptation.
  • For multi-lingual NLP, Zhao et al. [251] aimed to enhance training efficiency in multilingual NLP by adapting pre-trained FedFM through prompt tuning. FL-MetaSend [252] used knowledge distillation to selectively transfer global knowledge based on entropy measures, improving generalization across various domains. Liu et al. [253] explored different clustering methods to group adapter parameters, reducing the negative impact of data diversity. They found that language family-based clustering significantly outperforms other strategies. Weller et al. [254] demonstrated that fine-tuning pre-trained FedFM can achieve performance comparable to centralized fine-tuning in NLP settings. FedKC [255] employs k-means clustering on client data to generate representative knowledge, specifically in the form of clustered data centroids. These centroids are subsequently shared among clients for local training, enhancing data richness and addressing heterogeneity issues.
VIII Lessons Learned on FedFM
This section summarizes the key insights and takeaways from Sections IV- VII-A, providing readers with a clear understanding of the practical implications and broader context of the findings from this survey. In addition, we have connected these lessons to the topics discussed in Section IX to form a cohesive narrative.
Section IV focuses on efficient training and aggregation methods in FedFM. In FedFM, the heterogeneity of client devices in terms of computational resources and data distributions necessitates sophisticated customization of model architectures and training protocols. Techniques such as PEFT and Prompt Tuning are pivotal for optimizing performance and efficiency. PEFT methods, including adapter tuning, BitFit and LoRA, allow selective fine-tuning of a small subset of parameters, significantly reducing computational and communication overhead while maintaining high model performance. These methods are crucial for adapting large models to resource-constrained environments by selectively updating the most relevant parameters. Prompt tuning involves incorporating task-specific prompts into the input data, providing an efficient means of adapting pre-trained models to new tasks with minimal additional training. Robust and efficient aggregation methods, such as hierarchical aggregation, are also essential to reduce communication costs by aggregating updates at multiple levels. The key lesson here is the necessity of developing adaptive and efficient training and aggregation strategies that can cater to the heterogeneous environments typical of FedFM deployments, ensuring robust and scalable model training.
Section V focuses on the critical areas of robustness, privacy, and intellectual property for FedFM. Ensuring robustness and privacy in FedFM is critical due to the sensitivity of the data involved and the potential for malicious attacks. Byzantine robustness addresses the challenge of dealing with malicious or faulty clients that can disrupt the learning process. Techniques such as robust aggregation methods, including geometric median and trimmed mean, help in mitigating the impact of Byzantine clients by aggregating updates in a way that reduces the influence of outliers. Privacy preservation is equally crucial, with techniques like DP and secure aggregation playing a vital role. Protecting intellectual property is vital to securing the ownership of FedFM from unauthorized exploitation, including model theft. The lesson learned here is the critical importance of developing methods that balance robustness and privacy with performance, meanwhile protecting intellectual property of FedFM. Continuous advancements in cryptographic techniques and robust aggregation algorithms are essential for enhancing the security and resilience of FedFM systems against adversarial threats.
Section VI focuses on the incentive mechanism and contribution evaluation for FedFM. The successful deployment of FedFM relies heavily on the active participation of diverse stakeholders, necessitating effective incentive mechanisms. These mechanisms must ensure fairness, transparency, and security, motivating data owners to contribute their resources while deterring malicious behavior. Contract theory-based approaches can formalize participation terms, ensuring stakeholders are fairly compensated. Game theory-based methods model interactions between participants as strategic games, designing mechanisms that promote cooperative behavior and optimize collective outcomes. Auction-based methods provide a dynamic approach to incentivization, allowing participants to bid for the opportunity to contribute or receive rewards based on their resources and data quality. A critical takeaway from this section is that incentive mechanisms must be tailored to the specific needs and behaviors of FedFM participants. Combining multiple incentive strategies can address diverse motivations, ensuring sustained and meaningful participation, which is crucial for the success of FedFM.
Section VIII focuses on the implementation and applications of FedFM. The implementation of FedFM requires a comprehensive framework that integrates various components essential for effective deployment. This section discusses a structured approach to implementing FedFM, encompassing data management strategies, model training protocols, and efficient communication mechanisms. The framework emphasizes modularity, allowing for seamless integration of new methods and technologies as they evolve. Best practices in data preprocessing, feature extraction, and model optimization are incorporated to ensure robust and scalable implementations. In addition, this section highlights the transformative potential of FedFM across various applications, including NLP, speech recognition, recommendation systems, and healthcare. The lesson learned is that leveraging the comprehensive structure provided by the FedFM framework, along with its scalability and adaptability, can significantly enhance the effectiveness and reach of FL implementations, ensuring they meet the specific requirements of various real-world applications.
Figure 14: Challenges and promising directions for FedFM.
In summary, the lessons learned from Section VIII provide a foundation for the envisioned future directions and open challenges discussed in Section IX. By understanding the practical implications and challenges identified in our survey, researchers and practitioners can better develop key areas such as advanced customization, enhanced privacy mechanisms, comprehensive incentive strategies, and scalability solutions. Future research should focus on developing more adaptive strategies that can dynamically adjust to diverse client environments, robust privacy-preserving techniques that minimize performance trade-offs, and comprehensive incentive mechanisms that integrate multiple approaches to effectively motivate diverse participants. Additionally, exploring scalable and adaptable solutions will be crucial for meeting the evolving demands of real-world applications and advancing the field of FedFM.
IX Challenges and Promising Directions
@@
IX-A Memory and Computing Challenges facing FedFM
@@
Multi-modal GPT-4o: The large scale of FMs, often involving billions of parameters, results in significant memory demands. This is particularly challenging in FL environments, where client devices may have limited memory capacities. The key challenge is developing strategies to accommodate these large models efficiently within the memory constraints of diverse client devices. For multi-modal GPT-4o Model, it integrates text, audio, and vision capabilities in real-time, promising to transform human-machine interaction[256, 257]. Further FedFM research can focus on seamless integration and processing of multi-modal data from diverse sources while preserving privacy. Enhancing cross-modal learning in federated settings enables models to leverage insights from different data types effectively. In addition, optimizing the performance of multi-modal models like GPT-4o in FL frameworks is vital, particularly for communication efficiency and computational demands.
In FedFM, the demand for high computational resources on FL client devices poses a persistent challenge. Despite strategies to mitigate this, it remains an open issue. Approaches such as model splitting [258] and model compression [259] have been proposed. In addition, efforts are made to develop lighter models and more efficient training algorithms [260] that are less taxing on device resources. Yet, these solutions only offer a partial remedy. The research on federated optimization [261, 33] indicate that finding the right balance between maintaining model performance and managing resource constraints is an ongoing struggle. Although these advances mark progress, ensuring that FedFM is adaptable for devices with diverse capabilities continues to be a significant open research question.
@@
IX-B Communication Challenges facing FedFM
@@
Security Issues: In FedFM supported by advanced network technologies like 5G and beyond (B5G), security issues pose a significant challenge. The diversity in computing and communication capabilities across different participants in the network, stemming from variations in hardware (e.g., CPU, GPU), network connections (e.g., 4G, 5G, B5G, WiFi) and energy resources (e.g., battery, charged), leads to system heterogeneity  [287, 288, 289, 290]. This diversity can introduce inconsistencies and vulnerabilities in FedFM [291]. Moreover, the presence of unreliable devices within the network might lead to Byzantine failures as highlighted in [292, 293]. These failures refer to scenarios where certain nodes in the network act in a faulty or malicious manner, further complicating the security landscape of FedFM. The varying levels of security across different devices exacerbate the difficulty in defending against attacks and ensuring system reliability. Despite ongoing efforts to enhance FL security [23], these inherent vulnerabilities in FedFM remain open.
@@
GPT Model Management: Currently, there are millions of new GPTs in the GPT Store. The GPT store concept envisions a vast repository of pre-trained GPT models, each specialized for different tasks, domains, or languages. FedFM can explore dynamic selection and personalization of the most suitable GPT models for specific participant needs. In addition, efficient FedFM techniques to aggregate updates from diverse GPT models while maintaining model performance and privacy are crucial. Optimal management of computational and data resources is also essential to support millions of GPT models in a federated environment, ensuring robust performance and low latency.
IX-C Promising Direction: Quantum Computing for FedFM
@@
The ability of quantum computing to efficiently perform certain types of sampling and approximation (e.g., Monte Carlo sampling) can be advantageous for many ML tasks. For example, [302] have demonstrated a framework for reservoir computing with nonlinear quantum reservoirs, showcasing the computational capabilities across classical and quantum regimes. Quantum systems enable more efficient data structure representation, which can expedite computation in NNs during inference. When fine-tuning FedFM for specific tasks, feature selection is a crucial process. Quantum computing can offer efficient solutions for this process, enhancing the performance of the resulting FMs [303]. Lastly, quantum parallelism can support simultaneous evaluation of different hyper-parameter settings. This capability is significant for the hyper-parameter tuning process, a time-consuming aspect of FM fine-tuning. Moreover, the capability of quantum computing for handling high-dimensional spaces can aid in optimizing the high-dimensional parameter space of FMs more efficiently.
@@
The integration of quantum and classical optimization steps in hybrid algorithms offers a balanced approach. By utilizing quantum processors for the more computationally demanding parts and classical processors for other parts, these hybrid models leverage the strengths of both quantum and classical computing. They can be particularly effective in overcoming current limitations of quantum computing, while still harnessing its advantages for optimization tasks [307]. Global optimization is a key factor in FedFM settings, where the objective is to achieve consensus among distributed models. Quantum algorithms can efficiently aggregate local updates and conduct global optimization. This can produce more effective and cohesive FMs in a collaborative learning environment. The incorporation of quantum optimization into FedFM represents a promising direction. By enhancing the efficiency of finding optimal solutions, these quantum-based techniques have the potential to significantly improve the performance and effectiveness of FedFM.
@@
IX-D Promising Direction: Model Evaluation of FedFM